# Mil2: Efficient Cloth Simulation Using Non-distance Barriers and Subspace Reuse

### v3

@@ Abstract.
Mil2 pushes the performance of high-resolution cloth simulation, making the simulation interactive (in milliseconds) for models with one million degrees of freedom (DOFs) while keeping every triangle untangled. The guarantee of being penetration-free is inspired by the interior-point method, which converts the inequality constraints to barrier potentials. Nevertheless, we propose a major overhaul of this modality by defining a novel and simple barrier formulation which does not depend on the distance between mesh primitives. Such a non-distance barrier model allows a new way to integrate collision detection into the simulation pipeline. Another contributor to the performance boost comes from the so-called subspace reuse strategy. This is based on the observation that low-frequency strain vibrations are near orthogonal to the deformation induced by collisions or self-collisions, often of high frequency. Subspace reuse then takes care of low-frequency residuals, while high-frequency residuals can also be effectively smoothed by GPU-based iterative solvers. We show that our method outperforms existing fast cloth simulators by nearly one order while keeping the entire simulation penetration-free and producing high-equality animations of high-resolution models.
@@
Figure 1. Fashion show.  We present a cloth simulation framework that can be conveniently parallelized on the GPU. We name this new pipeline Mil2 as it is able to simulate high-resolution cloth meshes (e.g., around one million degrees of freedom) at an interactive rate (in milliseconds). Mil2 offers new insights into cloth and deformable body simulation. With a non-distance-based barrier formulation, we save a lot of computations of traditional CCD with a partial CCD procedure, which is much less expensive. The subspace reuse strategy relaxes the low-frequency errors effectively at the cost of single-digit milliseconds. Mil2 also features a residual forwarding trick to alleviate the damping issues generated by small-step line search filtering. In the teaser, we show an animated scene of the virtual fashion show. The model, dressed in a soft and light midi skirt, walks to the front and then turns around. These series of movements cause complex fabric dynamics, vividly showcasing the design concept of the garment. The garment is of high resolution and has $340$K vertices. The corresponding simulation involves over one million unknowns, and detailed local wrinkles can be well perceived. With a time step of $1/200$ sec, Mil2 runs at $5.4$ FPS.
@@ 1. Introduction
Being one of the most researched problems in computer graphics, cloth animation brings the simulated world to life in a vivid way, endows virtual characters with an infinite array of new appearances, and allows artists to unleash their talents and inspiration on the triangular mesh. The primary challenge for today’s cloth simulation arises from the irreconcilability between the desired visual quality and the limited computing resources – it is often the case that the actual time budget allocated for the simulator is strictly capped e.g., in interactive design or games. Cloths and fabrics demonstrate intricate dynamics under collisions and contacts, yielding captivating fine deformations of wrinkles and folds. To faithfully capture those effects, a high-resolution mesh is needed, and the increased number of degrees of freedom (DOFs) further stresses the simulation performance. This paper presents a GPU-based cloth simulation framework that is one or even two orders faster than the state-of-the-art GPU simulation algorithms for high-resolution cloth models. We name our method Mil2 suggesting it enables an interactive frame rate (in milliseconds) for simulation scenes of a million DOFs.
@@
Efficient processing collision and self-collision of cloth is the pivotal concern for a high-performance cloth simulator. We are inspired by the recent success of the barrier method a.k.a. incremental potential contact or IPC (Li et al., 2020, 2021; Lan et al., 2022b), and choose it as the primary modality for contact handling. IPC introduces a nonlinear repulsion between a pair of colliding or near-colliding primitives, which becomes infinitely strong if they get closer to each other. While this method has been proven robust, it requires repetitive CCDs (continuous collision detection) to calculate the distance between primitives in proximity – this is costly for detailed meshes since all the triangles may be in contact in cloth animations. As a result, the most dominant computation in many cloth simulators becomes the CCD processing. To address this issue, Mil2 uses a novel barrier formulation, which does not depend on the actual distance between primitives. Instead, the barrier energy is defined based on the virtual life span of a collision event – the “longer” a constraint/contact persists during the nonlinear solve, the stiffer it should be. Because of that, most CCD procedures invoked in the simulation can be substantially simplified.
@@
On the solve side, we observe that existing GPU-based iterative solvers are less effective for smoothing low-frequency errors. In theory, this drawback could be remedied with model reduction techniques, which project the system into a prescribed kinematic subspace. The real difficulty comes from the nonlinearity of the system matrix, which varies under different cloth poses, and constructing a subspace for each pose is not an option. To this end, Mil2 employs a subspace reuse scheme that leverages a low-frequency rest-shape subspace for different deformed poses. This strategy well synergizes with the projective dynamics (PD) framework (Bouaziz et al., 2014) because the geometric nonlinearity of cloth dynamics is taken care of in the local step, and the low-frequency subspace tends to be less sensitive to high-frequency deformations. We carefully exploit the structure of the global matrix in PD so that the expensive subspace projection of the full global matrix can be pre-computed. Our experiments show that subspace reuse cut the follow-up aggregated Jacobi iteration by $70\%$ on average.
@@
Mil2 is also equipped with a residual forwarding scheme. As time-critical applications allocate a limited time budget for the simulator, early termination may lead to visual artifacts like over-stiffening and damping due to the CCD-based line search filtering. Residual forward estimates a ghost external force as the residual force inherited from the previous step to relieve this issue.
@@
In a nutshell, Mil2 is a parallel simulator integrating several novel features to address the core challenges in cloth simulation. By deeply optimizing the pipeline, Mil2 further pushes the quality and efficiency of cloth simulation. Thanks to the barrier method, it is guaranteed that the resulting cloth poses are free of inter-penetration. Meanwhile, the combination of reused subspace and aggregated Jacobi iterations makes the solver effective for both low- and high-frequency deformations. More importantly, most calculations are friendly for GPU or any parallel computation platforms. Our experiments show that Mil2 delivers high-quality animation results while being one order faster than existing methods. Just as suggested by its naming, Mil2 enables an interactive frame rate (in milliseconds) for high-resolution cloth simulations (of a million DOFs).
@@ 2. Related Work
The cloth is regarded as a codimensional object (Terzopoulos et al., 1987), whose thickness is much smaller than its other two dimensions. A common practice is to discretize its geometry with a mass-spring network (Choi and Ko, 2002; Liu et al., 2013) or a triangle mesh (Baraff and Witkin, 1998; Volino et al., 2009). Early techniques use explicit integration with small time steps (Provot et al., 1995). The stability is improved by switching to the implicit integration (Baraff and Witkin, 1998), at the cost of assembling and solving the resulting linearized systems. Cloth fabric is less extensible, showing strong resistance to stretching. The discrepancy between its stretching and bending behavior induces extra numerical difficulty. Straining limiting is a simple and effective approach to mitigate this challenge (Provot et al., 1995; Thomaszewski et al., 2009; Wang et al., 2010). On the other hand, being unstretchable also inspires a simplified quadratic bending model (Bergou et al., 2006). Cloth is more than an isotropic hyperelastic continuum. Kim (2020) reveals the underlying connection between the Baraff-Witkin model (Baraff and Witkin, 1998) and anisotropic FEM. One can finetune the strain-stress relation to obtain an accurate material model i.e., see (Volino et al., 2009). Data-driven methods have also been used for this purpose (Wang et al., 2011; Sperl et al., 2022; Feng et al., 2022).
Collision processing for cloth
Due to the high surface-to-volume ratio, collision handling has always been an indispensable part of cloth simulation. Based on CCD, Provot (1997) and Bridson et al. (2002) suggest a post-simulation step grouping penetrating vertices into impact zones. Huh et al. (2001) decomposes the impact zone into smaller colliding clusters based on their positions on the original mesh. Harmon et al. (2008) relaxes impact zones with an inelastic projection, allowing relative tangential movement. CCD is known to be expensive – it solves a cubic function for the first time of impact (TOI). The analytic formulation of cubic function is available but error-prone (Wang et al., 2021). Numerical root finding is therefore preferred (Wang, 2014; Yuksel, 2022), which is clearly more costly.
In time-sensitive cases, discrete collision detection (DCD) may be used as a cheap alternative. For collisions between cloth and volumetric colliders (e.g., the human body), the penalty force based on projection (Guan et al., 2012) or signed distance field (SDF) (Chen et al., 2013) are popular choices. Buffet et al. (2019) extends the implicit field of volumetric objects to open surfaces to resolve the inter-penetration of multiple-layer cloth. For collisions between cloths, Baraff et al. (2003) and Wicke et al. (2006) employ an untangling method that applies repulsion forces to minimize the colliding region. Volino and Magnenat-Thalmann (2006) and Ye and Zhao (2012) minimize the length of the collision contour of the colliding region.
@@
Recently, an interior-point-based algorithm called Incremental Potential Contact (IPC) (Li et al., 2020) has been proposed, which guarantees the simulation to be free of intersections. It has then been generalized to the simulation of cloths/thin shells (Li et al., 2021), rigid/stiff bodies (Ferguson et al., 2021; Lan et al., 2022a), and curved meshes (Ferguson et al., 2023). IPC is highly time-consuming as a CCD is needed at each nonlinear iteration to ensure the simulation results within the feasible region. Lan et al. (2022b) dynamically approximates the logarithm barrier with an increasingly stronger quadratic function so that the simulation fits the PD framework (Bouaziz et al., 2014). Lan et al. (2023) decompose the global collision configuration into local stencils and solve them in parallel. To reduce the cost of CCD, Wu et al. (2020) uses point-point distance constraints between triangle pairs to avoid intersection.
@@
In addition to collision, another computational bottleneck is the (nonlinear) system solve due to the use of implicit integration. A widely used strategy is to convert the force equilibrium (i.e., strong form) to the variational form (i.e., weak form) (Kharevych et al., 2006; Gast et al., 2015; Li et al., 2019). Doing so offers new perspectives to the simulation in the light of optimization a.k.a. constraint-based methods. One can locally and inexactly solve those constraints via the so-called constraint projection (Goldenthal et al., 2007). For instance, position-based dynamics (Müller et al., 2007; Macklin et al., 2016) uses per-vertex constraint projection making the simulation matrix-free. Projective dynamics (PD) (Bouaziz et al., 2014) presents a global and local alternation scheme to solve the nonlinear dynamic system. PD quickly becomes a popular simulation modality because its local projections are trivially parallelizable. Instead of solving its global system exactly e.g., using Cholesky factorization, iterative linear solvers can be used, such as Jacobi (Wang, 2015; Lan et al., 2022b), Gauss-Seidel (Fratarcangeli et al., 2016) and preconditional conjugate gradient (PCG) (Tang et al., 2013), which bring significant speedups on the GPU. For more general and nonlinear models, sophisticated GPU algorithms are needed to decouple unknown DOFs to obtain the global solution (Wang and Yang, 2016; Tang et al., 2018; Wang et al., 2023; Lan et al., 2023). In cloth simulation, our human perception is sensitive to the richness of local details. Given limited computation time, we often opt for incorporating more DOFs over more complicated materials for improved visual plausibility.
@@
Model reduction is an acceleration technique that reduces the simulation cost. It uses a set of reduced coordinates to pre-parameterize the simulation in the subspace and lowers the total number of unknown DOFs. Linear modal analysis offers the optimal subspace approximate around the rest shape (Pentland and Williams, 1989; Choi and Ko, 2005; O’Brien et al., 2003). Large and rotational deformations are less intuitive. For StVK models, one can pre-compute the coefficients of the reduced Hessian and the internal force (Sifakis and Barbic, 2012). An et al. (2008) present a data-driven approach to allow the integration to be performed only at a few sample elements (cubature). It is also possible to combine modal analysis with stiffness warping (Müller et al., 2002) at per-vertex local frames (Choi and Ko, 2005). However, doing so suffers from ghost forces when simulating free-floating objects. Another collection of contributions builds data-driven reduced models. For instance, Kim and James (2009) use recent simulation results to construct the subspace at the simulation runtime. Shen et al. (2021) and Fulton et al. (2019) use an autoencoder to encode the fullspace DOFs using the latent representation. Reduced models need expensive pre-computations such as computing the generalized eigendecomposition or training a deep network. While some efforts have been made to speed up such pre-computations (Yang et al., 2015), any changes to the simulation setup e.g., the boundary condition, nullify the pre-computation. On the contrary, Mil2 employs a novel subspace reuse method – the subspace constructed at the rest shape remains effective, and one only needs incremental computations for the subspace integration.
@@
Multigrid method (Trottenberg et al., 2001) is another common technique to boost simulation efficiency when a large number of DOFs are present. It was originally proposed to solve Poisson-like equations abounded in fluid simulation (McAdams et al., 2010; Molemaker et al., 2008). For deformable/cloth simulation, one needs to represent the dynamics at different levels. The geometric multigrid (GMG) (Georgii and Westermann, 2006) approaches this by generating spatial discretization (e.g., meshes or grids) of different resolutions. Xian et al. (2019) further simplifies this process by sampling points from the finest grid to form coarser grids. The algebraic multigrid (AMG), on the other hand, approaches this by generating a subspace of the fine dynamics, which shares a similar nature of model reduction. For example, Li et al. (2023) utilizes a B-spline subspace, and Tamstorf et al. (2015) built the subspace by QR decomposition on near-kernel components. Though most works treat the multigrid method as a linear solver, Wang et al. (2018) integrates multigrid into a nonlinear optimization process, which updates the residue and system matrix periodically. Mil2 can be understood as a two-level multigrid for the global step solve. The reused subspace solve eliminates the low-frequency errors, leaving the high-frequency to the aggregated Jacobi iterations.

### v4

# Efficient GPU Cloth Simulation with Non-distance Barriers and Subspace Reuse

@@ Abstract.
This paper pushes the performance of cloth simulation, making the simulation interactive even for high-resolution garment models while keeping every triangle untangled. The penetration-free guarantee is inspired by the interior point method, which converts the inequality constraints to barrier potentials. We propose a major overhaul of this modality within the projective dynamics framework by leveraging an adaptive weighting mechanism inspired by barrier formulation. This approach does not depend on the distance between mesh primitives, but on the virtual life span of a collision event and thus keeps all the vertices within feasible region. Such a non-distance barrier model allows a new way to integrate collision resolution into the simulation pipeline. Another contributor to the performance boost comes from the subspace reuse strategy. This is based on the observation that low-frequency strain propagation is near orthogonal to the deformation induced by collisions or self-collisions, often of high frequency. Subspace reuse then takes care of low-frequency residuals, while high-frequency residuals can also be effectively smoothed by GPU-based iterative solvers. We show that our method outperforms existing fast cloth simulators by at least one order while producing high-quality animations of high-resolution models.
@@
ccs: Computing methodologies Physical simulation
Figure 1. Fashion show.  We present a new GPU-based cloth simulation framework with projective dynamics. Our method is able to simulate high-resolution cloth meshes at an interactive rate. With a non-distance-based barrier formulation, we can replace a large portion of traditional CCDs with the partial CCD procedure, which is much less expensive. The subspace reuse strategy relaxes the low-frequency errors effectively at the cost of single-digit milliseconds. Our method also features a residual forwarding trick to alleviate the damping issues generated by early termination and small-step line search filtering. In the teaser, we show an animated scene of the virtual fashion show. The model, dressed in a soft and light midi skirt, walks to the front and then turns around. These series of movements cause complex fabric dynamics, vividly showcasing the design concept of the garment. The garment is of high resolution and has $340$K vertices. The corresponding simulation involves over one million unknowns, and detailed local wrinkles can be well perceived. With a time step of $\Delta t=1/200$ sec, the simulation runs at $4.8$ FPS. Please refer to the supplementary video for the corresponding animations.
@@ 1. Introduction
Cloth animation brings the simulated world to life in a vivid way, endows virtual characters with an infinite array of new appearances, and allows artists to lay their talents and inspirations on the triangular mesh. The primary challenge for today’s cloth simulation arises from the irreconcilability between the desired visual quality and the limited computing resources – it is often the case that the actual time budget allocated for the simulator is strictly capped e.g., in interactive design or games. Cloths and fabrics demonstrate intricate dynamics under collisions and contacts, yielding captivating fine deformations of wrinkles and folds. To faithfully capture those effects, a high-resolution mesh is preferred, and the increased number of degrees of freedom (DOFs) further stresses the simulation performance. This paper presents a GPU-based cloth simulation framework which is one or even two orders faster than the state-of-the-art GPU simulation algorithms for high-resolution cloth models.
@@
Efficient processing of collision and self-collision of cloth is the pivotal concern for a high-performance cloth simulator. Our method is inspired by the recent success of incremental potential contact or IPC (Li et al., 2020, 2021a). When combined with CCD line search filtering and projection-Newton solver, IPC offers a non-penetration guarantee throughout the simulation process. It introduces a nonlinear repulsion between a pair of colliding or near-colliding primitives, which becomes infinitely strong if they get closer to each other. While this method has been proven robust, it requires repetitive CCDs (continuous collision detection) to calculate the distance between primitives in proximity – this is costly for detailed meshes since all the triangles may be in contact. As a result, a dominant computation in many cloth simulators becomes the CCD processing. To address this issue, we design a novel barrier mechanism, which adaptively re-weights the collision constraints when the collision event exists over a continuous period of iterations, and does not depend on the actual distance between primitives. Because of this, most CCD procedures invoked in the simulation can be substantially simplified.
@@
On the solve side, we observe that existing GPU-based iterative solvers are less effective for smoothing low-frequency errors. In theory, this drawback could be remedied with model reduction or multigrid techniques, which project the system into a prescribed kinematic subspace. The difficulty comes from the nonlinearity of the system matrix, which varies under changing cloth poses, and constructing a subspace for each pose is not an option. To this end, we design a subspace reuse scheme that leverages a low-frequency rest-shape subspace for different deformed poses. This strategy well synergizes with the projective dynamics (PD) framework (Bouaziz et al., 2014) because the geometric nonlinearity of cloth dynamics is taken care of in the local step, and the low-frequency subspace tends to be less sensitive to high-frequency deformations. We carefully exploit the structure of the global matrix in PD so that the expensive subspace projection of the full global matrix can be pre-computed. Our experiments show that subspace reuse cut the follow-up aggregated Jacobi iteration by $70\%$ on average.
@@
Our algorithm is also equipped with a residual forwarding scheme. As time-critical applications allocate a limited time budget for the simulator, early termination may lead to visual artifacts like over-stiffening and damping due to the CCD-based line search filtering. Residual forward estimates a ghost external force as the residual force inherited from the previous step to relieve this issue.
@@
In a nutshell, this paper proposes a simulator integrating several novel features to address the core challenges in cloth simulation. By deeply optimizing the pipeline, our method further pushes the quality and efficiency of the simulation. It is guaranteed that the resulting cloth poses are free of inter-penetration. Meanwhile, the combination of reused subspace and aggregated Jacobi iterations makes the solver effective for both low- and high-frequency deformations. More importantly, most calculations are friendly for GPU or any parallel computation platforms. The experiments show that our method delivers high-quality animation results while being one order faster than existing methods.
@@ 2. Related Work
A common practice for cloth simulation is to discretize its geometry with a mass-spring network (Choi and Ko, 2002; Liu et al., 2013) or a triangle mesh (Terzopoulos et al., 1987; Baraff and Witkin, 1998; Volino et al., 2009). Early techniques use explicit integration with small time steps (Provot et al., 1995). The stability is improved by switching to the implicit integration (Baraff and Witkin, 1998), at the cost of assembling and solving the resulting linearized systems. Cloth fabric is less extensible, showing strong resistance to stretching. The discrepancy between its stretching and bending behavior induces extra numerical difficulty. Strain limiting is a simple and effective approach to mitigate this challenge (Provot et al., 1995; Thomaszewski et al., 2009; Wang et al., 2010). On the other hand, being unstretchable also inspires a simplified quadratic bending model (Bergou et al., 2006). Cloth is more than an isotropic hyperelastic continuum. Kim (2020) reveals the underlying connection between the Baraff-Witkin model (Baraff and Witkin, 1998) and anisotropic finite element method (FEM). One can finetune the strain-stress relation to obtain an accurate material model i.e., see (Volino et al., 2009). Data-driven methods have also been used for this purpose (Wang et al., 2011; Sperl et al., 2022; Feng et al., 2022).
Collision processing
Collision handling has always been an integral part of cloth simulation. Discrete collision detection (DCD) is an efficient method to identify the list of colliding primitives. As the name suggests, DCD detects the inter-penetration at a specific time instance, and it fails to capture all the collision events for models with thin geometries (such as garments) or fast-moving objects. On the other hand, CCD offers a more robust way to detect inter-penetration. Assuming the trajectory is linear within a time interval, CCD calculates the first time of impact (TOI) between two primitives. For triangularized surfaces, CCD is often modeled as the root finding of cubic equations. It is error-prone (Wang et al., 2021), and robust numerical procedures are preferred (Wang, 2014; Yuksel, 2022). The brute-force collision detection among all the primitive pairs is prohibitive in general. A commonly adopted method is to use some bounding volume hierarchy (BVH) (Langetepe and Zachmann, 2006) to avoid excessive triangle-triangle intersection tests, a.k.a collision culling. Various BVH types have been explored such as AABB (Bergen, 1997), OBB (Gottschalk et al., 1996), sphere (Hubbard, 1995; James and Pai, 2004), Boxtree (Zachmann, 2002), spherical shell (Krishnan et al., 1998) and so on.
To resolve the intersected triangles, the penalty method has been a popular choice for its simplicity (Guan et al., 2012). Buffet et al. (2019) extend the implicit field of volumetric objects to open surfaces to resolve the inter-penetration of multiple-layer cloth. Baraff et al. (2003) and Wicke et al. (2006) employ an untangling method that applies repulsion forces to minimize the colliding region for cloth-cloth collisions. Volino and Magnenat-Thalmann (2006) and Ye and Zhao (2012) minimize the length of the collision contour of the colliding region. Instead of solving the collision and cloth dynamics in a two-way coupled manner, Provot (1997) and Bridson et al. (2002) suggest a post-simulation step grouping penetrating vertices into impact zones. Huh et al. (2001) decompose the impact zone into smaller colliding clusters based on their positions on the original mesh. Harmon et al. (2008) relax impact zones with an inelastic projection, allowing relative tangential movement.
@@
Recently, an interior-point-based algorithm called incremental potential contact (IPC) (Li et al., 2020) has been proposed, which guarantees the simulation to be free of intersections. It has then been generalized to the simulation of cloth/thin shells (Li et al., 2021a), rigid/stiff bodies (Ferguson et al., 2021; Lan et al., 2022a), and curved meshes (Ferguson et al., 2023). IPC is highly time-consuming as a CCD is needed at each nonlinear iteration to ensure the simulation results stay within the feasible region. Lan et al. (2022b) approximate the logarithm barrier with an increasingly stronger quadratic function so that the simulation fits the PD framework (Bouaziz et al., 2014). Lan et al. (2023) decompose the global collision configuration into local stencils. To reduce the cost of CCD, Wu et al. (2020) use point-point distance constraints between triangle pairs to avoid intersections. Another relevant prior art is from Ly et al. (2020). This method is also based on PD, and it incorporates Signorini-Coulomb law (Daviet et al., 2011; Brogliato and Brogliato, 1999) using constraint projections in a semi-implicit way. With a matrix splitting scheme, it handles frictional contact robustly and efficiently. Our method is orthogonal to (Ly et al., 2020) in the sense that we aim to improve the performance for contact resolution using a non-distance barrier-like formulation to adaptively adjust the collision weight instead of resorting to complementarity programming (Moreau, 1988). Our subspace reuse technique may also be used in (Ly et al., 2020).
@@
In addition to collision, another computational bottleneck is the (nonlinear) system solve due to the use of implicit integration. A widely used strategy is to convert the force equilibrium (i.e., strong form) to the variational form (i.e., weak form) (Kharevych et al., 2006; Gast et al., 2015; Li et al., 2019). Doing so offers new perspectives to the simulation in the light of optimization such as constraint-based methods. One can locally and inexactly solve those constraints via the the constraint projection (Goldenthal et al., 2007), which highlights the potential for parallelization. For instance, position-based dynamics (Müller et al., 2007; Macklin et al., 2016) uses per-vertex constraint projection making the simulation matrix-free. Projective dynamics (PD) (Bouaziz et al., 2014) presents a global and local alternation scheme to solve the nonlinear dynamic system. PD quickly became a popular simulation modality because its local projections are trivially parallelizable. Instead of solving its global system exactly e.g., using Cholesky factorization, iterative linear solvers can be used, such as Jacobi (Wang, 2015; Lan et al., 2022b), Gauss-Seidel (Fratarcangeli et al., 2016) and preconditional conjugate gradient (PCG) (Tang et al., 2013), which bring significant speedups on the GPU. For more general and nonlinear models, sophisticated GPU algorithms are needed to decouple unknown DOFs to obtain the global solution (Wang and Yang, 2016; Tang et al., 2018; Wang et al., 2023; Lan et al., 2023).
@@
Model reduction is an acceleration technique that reduces the simulation cost. It uses a set of reduced coordinates to pre-parameterize the simulation in the subspace and lowers the total number of unknown DOFs. Linear modal analysis offers the optimal subspace approximate around the rest shape (Pentland and Williams, 1989; Choi and Ko, 2005; O’Brien et al., 2003). Large and rotational deformations are less intuitive. For StVK models, one can pre-compute the coefficients of the reduced Hessian and the internal force (Sifakis and Barbic, 2012). It is also possible to combine modal analysis with stiffness warping (Müller et al., 2002) at per-vertex local frames (Choi and Ko, 2005). However, doing so suffers from ghost forces when simulating free-floating objects. Another collection of contributions builds data-driven reduced models. For instance, Kim and James (2009) use recent simulation results to construct the subspace at the simulation runtime. Shen et al. (2021) and Fulton et al. (2019) use an autoencoder to encode the fullspace DOFs using the latent representation.
Model reduction can also be coupled with PD. For instance, Brandt et al. (2018) design a reduced model for both local projection and global solve. It is highly efficient when the global matrix is constant (at the cost of missing some local deformations). When the global matrix varies due to collision and contact events or mesh’s topology changes, efficiently estimating the updated subspace matrix becomes a challenge. A strategy is to use Woodbury formulation to leverage the pre-factorized global matrix for the rest shape and low-rank matrix update as in (Modi et al., 2021; Li et al., 2021b). Apart from those existing methods, we employ a novel subspace reuse method – the subspace constructed at the rest shape is used for relax low-frequency errors. This strategy requires minimum computational cost, while the high-frequency is left for GPU-based iterative solvers.
@@
Multigrid method (Trottenberg et al., 2001) is another common technique to boost simulation efficiency when a large number of DOFs is present. It was originally proposed to solve Poisson-like equations abounded in fluid simulation (McAdams et al., 2010; Molemaker et al., 2008). For deformable/cloth simulation, one needs to represent the dynamics at different levels. The geometric multigrid (GMG) (Georgii and Westermann, 2006) approaches this by generating spatial discretization (e.g., meshes or grids) of different resolutions. Xian et al. (2019) further simplifies this process by sampling points from the finest grid to form coarser grids. The algebraic multigrid (AMG), on the other hand, approaches this by generating a subspace of the fine dynamics, which shares a similar nature of model reduction. For example, Li et al. (2023) utilize a B-spline subspace, and Tamstorf et al. (2015) built the subspace by QR decomposition on near-kernel components. Wang et al. (2018) integrate multigrid into a nonlinear optimization process, which updates the residue and system matrix periodically. Our method can be understood as a two-level multigrid for the global step solve. The reused subspace solve eliminates the low-frequency errors, leaving the high-frequency to the aggregated Jacobi iterations.